# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v1

@@ Abstract
The integration of tools in LLM-based agents overcame the difficulties of standalone LLMs and traditional agents’ limited capabilities. However, the conjunction of these technologies and the proposed enhancements in several state-of-the-art works followed a non-unified software architecture resulting in a lack of modularity. Indeed, they focused mainly on functionalities and overlooked the definition of the component’s boundaries within the agent. This caused terminological and architectural ambiguities between researchers which we addressed in this paper by proposing a unified framework that establishes a clear foundation for LLM-based agents’ development from both functional and software architectural perspectives. Our framework, LLM-Agent-UMF (LLM-based Agent Unified Modeling Framework), clearly distinguishes between the different components of an agent, setting LLMs, and tools apart from a newly introduced element: the core-agent, playing the role of the central coordinator of the agent which comprises five modules: planning, memory, profile, action, and security—the latter often neglected in previous works. Differences in the internal structure of core-agents led us to classify them into a taxonomy of passive and active types. Based on this, we proposed different multi-core agent architectures combining unique characteristics of various individual agents. For evaluation purposes, we applied this framework to a selection of state-of-the-art agents, thereby demonstrating its alignment with their functionalities and clarifying the overlooked architectural aspects. Moreover, we thoroughly assessed four of our proposed architectures by integrating distinctive agents into hybrid active/passive core-agents’ systems. This analysis provided clear insights into potential improvements and highlighted the challenges involved in the combination of specific agents.
@@
Keywords LLM-based agents  $\cdot$ software architecture  $\cdot$ modularity  $\cdot$ security  $\cdot$ classification  $\cdot$ taxonomy  $\cdot$ core-agent
@@ 1 Introduction
Large Language Models (LLMs) excel in tasks like language modeling, text generation, question answering, sentiment analysis, natural language understanding, and commonsense reasoning [1]. However, standalone LLMs lacks other skills such as accessing external knowledge, information retrieval, mathematical reasoning, code evaluation, and numerous others. These functional shortcomings can be managed by AI agents by leveraging external tools and human feedback. An autonomous agent is a system interacting with an environment, sensing it, and acting on it over time following a certain agenda [2]. There are different classes of agents but in this paper, we will be focusing on LLM-based agents which, by combining the capabilities of LLMs and autonomous agents, can achieve a broad range of tasks [1].
@@
In fact, LLM-based agents have now become dominant in the landscape of AI agents due to several reasons. The first is the previously discussed versatile set of capabilities inherited from LLMs. Secondly, LLMs are regarded as potential steppingstones towards Artificial General Intelligence (AGI), offering hope for the development of general AI agents that can adapt to diverse scenarios [3]. Finally, the shift towards more natural conversational interfaces powered by LLMs is transforming the way humans interact with agents, enabling seamless and intuitive interactions.
@@
Understanding the potential of these agents and improving them necessitates a deep understanding of their structure. In LLM-based agents, besides the LLM who handles reasoning, there are other components that oversee the execution of tasks, ensure the security of the agent, and handle its memory [3]. But merely pinpointing existing functionalities is insufficient for developers and researchers. To provide a comprehensive overview of these agents, survey [4] proposes a comprehensive framework for their construction, consisting of four main modules which will be discussed thoroughly in the next section. Although it provides valuable insights into the functionality of each module in an agent, it overlooks their delineation from a software architecture perspective. This perspective is crucial for developers to establish a common base architecture to build upon and for researchers to improve. Our analysis of the state-of-the-art LLM-based agents reveals common limitations: the complexity of implementation, resulting in unstructured and ambiguous software architecture; lack of modularity and composability, making components non-reusable by other agent-based solutions; and difficulty in maintainability and introducing improvements to existing agents. These issues stem from the absence of a clear framework that emphasizes the architectural perspective of the agent as much as its functionalities.
@@
Our proposed LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF), emphasizes a clear delineation of each component within an LLM-based agent, defining their interactions within specified boundaries from both architectural and functional perspectives. In addition to LLMs, we introduced a new unit within the agent: the core-agent, which we classify into two types: active and passive core-agents. This proposition adds clarity to the capabilities of each component and describes the dynamics of interactions between modules within the agent. As a result, we alleviate the complexity of the architecture and improve the reusability of the components, promoting a shift from multi-agent systems to multi-core agents.
@@
In the evaluation section, we highlighted the advantages offered by our framework, from resolving terminological ambiguities to the introduction of enhancing modules like the security module. To validate the reliability of our framework, we applied it to existing solutions and identified the modules within each agent. This approach enabled us to assess the feasibility and requirements for merging one agent with another, ultimately aiming to create an agent with fully enhanced capabilities.
@@
Compared with previous works, the 5 main contributions of this paper can be summarized as follows:
@@
    Introducing a new terminology, core-agent, as a structural sub-component in LLM-based agents to improve modularity and promote more effective and precise communication among researchers and contributors in the field of agents and LLM technologies.
@@
    Modelling the internal structure of a core-agent by adapting the framework suggested by [4] that was originally meant to describe the whole agent from an abstract functional perspective.
@@
    Improving our modeling framework by augmenting the core-agent with a security module and introducing new methods within other modules.
@@
    Classifying core-agents into active core-agents and passive core-agents, explaining their differences and similarities, and highlighting their unique advantages.
@@
    Finally, introducing various multi-core agent architectures emphasizing that the hybrid one-active-many-passive core-agent architecture is the optimal setup for LLM-based agents.
@@ 2 Related Work
We will start off this section by providing a comprehensive overview of the foundational concepts upon which our work is based. First, Tool-augmented LLMs are a major advancement in NLP that combine the language understanding and generation capabilities of LLMs with the ability to interface with external tools and APIs. For instance, TALM [5] introduces models which can leverage a wide range of functionalities, from information retrieval to task planning and execution.
@@
Figure 1: Framework proposed by survey [4] for LLM-based agents
@@
3 LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF)
@@ 3.1 Core-Agent: Keystone component of an LLM-based agent
Figure 2: The core-agent as the central component of LLM-based agents
@@
The LLM acts as a cerebral entity covering cognitive tasks such as natural language understanding and comprehensive text generation based on specific context. Indeed, LLMs can communicate with core-agents which in turn can interact with tools or data sources within the system. Considering that an agent can manifest as unimodal or multimodal [17], the latter can be replicated with different task-specific LLMs, as illustrated in Figure 2. This distinction fosters modularity within our LLM-based agent unified modeling framework, making them open for extension and addition of further domain-specific models without having to introduce changes to the overall system architecture, adhering more to the Open-Close principle (OCP) [18] [19] .
@@
In LLM-based agents, tools are important assets. They can take various forms such as supplementary systems, software applications or even physical devices that extend and enhance the capabilities of an agent. They span across a spectrum of complexity, ranging from basic API integrations to sophisticated auxiliary systems designed for specific tasks. In this context, tools can be considered external if they are independent systems which achieve a complete objective, or internal if they cooperate with the core-agent to achieve tasks in the scope of the agent goal.
@@ 3.2 Modeling the Core-Agent Internal Structure
The human brain exhibits remarkable modularity, with distinct regions and functionalities working in coordination to facilitate cognitive processes, decision-making, and behavior. Inspired by this organized design, our proposed framework aims to emulate the brain’s modular structure thought the incorporation of five internal modules: the planning module, the memory module, the profile module, the action module and the security module. By integrating the latter concepts, the core-agent serves as the central coordinator, controlling the interactions and information flow between different components, as elaborated in the previous section.
@@
To provide a robust comparison and underscore the novelty of our framework, it is instructive to examine it alongside existing approaches. A recent survey [4] introduced a framework for LLM-based agents, comprising four key modules: profile, memory, planning, and action modules. For a comprehensive overview of this framework, we direct readers to the “Related Work” section. While this survey’s proposed framework [4] treated the LLM-based agent as a whole system without distinguishing a core component from other entities, our adaptation, tailored specifically for the core-agent, necessitates modifying certain definitions and module structures to accommodate the separation between the LLM and the core-agent. Notably, we had to reposition the four key modules to operate under the core-agent within the framework alongside our newly introduced security-oriented module.
Figure 3: The internal structure of a core-agent
@@ 3.2.1 Planning Module
This section will detail the functionalities and characteristics of this module from four critical aspects: process, strategies, techniques, and feedback sources. These aspects derive from the clear separation we have established between the core-agent and the LLM, and between our system and external systems Figure 2.
@@
    This initial phase involves breaking down a complex task into simpler subtasks, thereby establishing a structured hierarchy of intermediate goals. The decomposition of complex tasks can adopt two primary approaches: In the non-iterative decomposition approach, the complex task is broken down into simple subtasks all at once. The planning module defines all subtasks, creating a complete task hierarchy in a single step. This method provides a comprehensive overview of the entire task structure upfront. However, the iterative decomposition approach involves a step-by-step breakdown of the task. In fact, the planning module first defines the initial subtask and goal to reach, establish a proper plan and generates the procedures to undertake. After performing the procedures and completing the plan, the output is considered, and the next subtask is defined. This process is repeated, with each new subtask being planned and executed, contributing to defining the next subtask to achieve. The key advantage of this method is its flexibility and ability to adapt the plan based on the outcomes of each subtask. An exemplary application of the iterative approach is the Decomposition-Alignment-Reasoning Agent (DARA) framework [22]. DARA [22] demonstrates how iterative decomposition can lead to more precise and context-aware planning, particularly for complex tasks with multiple interdependent subtasks.
@@
    This approach involves generating a singular path or sequence of procedures to achieve the goal, adhering to the plan step-by-step without exploring alternatives, thereby providing a straightforward, deterministic approach to planning. Chain of thought (CoT) [23] is an example that outlines such a strategy. Indeed, it uses sequential reasoning as it involves breaking down complex problems into multiple procedures, each built on the previous ones [23].
@@
    As a matter of fact, the Tree of Thoughts (ToT) [24] and Graph of Thoughts (GoT) [24] are two frameworks that utilize this multi-path approach. They both operate by leveraging an LLM as a thought generator to produce intermediate procedures, that are structured either as a hierarchical tree in case of ToT or as a more complex graph in case of GoT. However, the complexity of managing these structures cannot be solely handled by the LLM. It requires the integration of a specialized software component responsible for orchestrating the process, interacting with the LLM, and organizing the thoughts into the desired structure, whether a tree or a graph. This essential software component is identified as the planning module within the core-agent. The planning module further evaluates different paths within the generated structure and selects an optimal plan [21] based on its assessment.
@@
The planning module follows planning techniques as methodological approaches to form executable plans. These techniques are chosen based on criteria such as the complexity of the task, and the need for contextual comprehension. Our framework present two primary techniques:
@@
    Within our architectural framework, rule-based methodologies encompass what is commonly referred to in literature as symbolic planners [21]. These techniques proved to be valuable especially in contexts characterized by complex constraints, such as mathematical problem-solving or the generation of plans within highly problematic situations [21].
@@
Planning without feedback may pose several challenges as feedback plays a crucial role in optimizing the performance of the planning module within the core-agent. For instance, in iterative task decomposition, feedback has an influential impact on the next generated step and enhances the agent’s alignment with the user’s expectations.
To effectively address these challenges, the planning module relies on diverse feedback sources. As outlined in Figure 3, the core-agent engages with tools within its system boundaries, as well as entities outside its scope, such as external systems and humans. Consequently, interactions with these components can offer valuable feedback:
@@ 3.2.2 Memory Module
The memory module is responsible for the storage and retrieval of information pertinent to the core-agent’s activities, thereby enhancing the core-agent’s decision-making efficiency and task execution capabilities. In [26] and [4], the memory module in an LLM-based agent was approached from an abstract functional perspective, neglecting its analysis from a software architectural viewpoint. This led to some overlap between the memory module and the other modules defined in the framework suggested by [4]. Consequently, we propose a more comprehensive and well-defined presentation of this module based on three perspectives: Memory Structure, Memory Location, and Memory Format.
The framework presented in the survey [4] classified the memory structure into Unified Memory, intended to emulate human short-term memory via in-context learning, and Hybrid Memory, which represents both short-term and long-term memory functionalities. While these names aimed to differentiate types of agent systems, they deviate unnecessarily from the well-established terminology in the field. A more conventional and widely recognized categorization is introduced through the first perspective, Memory Structure, which includes short-term and long-term memory, aligning more with the human memory structure as established by [27].
As opposite to how it was defined in framework [4], our short-term memory definition diverges to focus more on the impact of the core-agent rather than on the LLM. From a content perspective, various sources of information contribute to short-term memory. As pointed out in paper [26], data can be derived from one trial of a given task or from previous trials of the same task, which is referred to as short-term memory. This data has a narrow scope that focuses on a specific task and primarily communicated to the LLM for the purpose of in-context learning and enhancing the LLM capability with more information related to the task at hand. Conversely, long-term memory refers to the ability to store and recall information over extended periods, beyond the scope of a specific task. This type of memory enables the core-agent to maintain coherence and context over prolonged interactions, learning and adapting from past experiences. Namely, MemoryBank stores all interactions with user in a large symbolic memory and processes experiences into high-level summaries to reflect upon future similar tasks [28].
@@
Beyond focusing on the scope of the data, inside trial and across trials, as stated in [26], we emphasize the location of the memory because it is more relevant from a software architecture point of view. Hence, we introduce the third perspective, Memory Location, which encompass two categories: Embedded Memory, internal to the core-agent boundaries, and Memory Extension, outside the core-agent boundaries, yet still within the boundaries of the agent system as depicted in Figure 3.

### v2

@@ Abstract
In an era where vast amounts of data are collected and processed from diverse sources, there is a growing demand to develop sophisticated AI systems capable of intelligently fusing and analyzing this information. To address these challenges, researchers have turned towards integrating tools into LLM-powered agents to enhance the overall information fusion process. However, the conjunction of these technologies and the proposed enhancements in several state-of-the-art works followed a non-unified software architecture resulting in a lack of modularity and terminological inconsistencies among researchers. To address these issues, we propose a novel LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF) that aims to establish a clear foundation for agent development from both functional and software architectural perspectives. Our framework clearly distinguishes between the different components of an LLM-based agent, setting LLMs, and tools apart from a new element, the core-agent, playing the role of the central coordinator of the agent. This pivotal entity comprises five modules: planning, memory, profile, action, and security—the latter often neglected in previous works. By classifying core-agents into passive and active types based on their authoritative natures, we propose various multi-core agent architectures that combine unique characteristics of distinctive agents to tackle complex tasks more efficiently. We evaluate our framework by applying it to thirteen state-of-the-art agents, thereby demonstrating its alignment with their functionalities and clarifying the overlooked architectural aspects. Moreover, we thoroughly assess five of our proposed architectures through the integration of existing agents into new hybrid active/passive core-agents architectures. This analysis provides clear insights into potential improvements and highlights challenges involved in combining specific agents.
@@
Keywords LLM-based agent  $\cdot$ software architecture  $\cdot$ modularity  $\cdot$ privacy  $\cdot$ security  $\cdot$ classification  $\cdot$ multi-core agent
@@ 1 Introduction
Large Language Models (LLMs) excel in tasks like language modeling, question answering, sentiment analysis, Natural Language Understanding (NLU), commonsense reasoning and knowledge fusion [1, 2]. However, standalone LLMs lacks other skills such as information retrieval, mathematical reasoning, code evaluation, and numerous others. These functional shortcomings can be managed by AI agents leveraging external tools, knowledge repositories and human feedback. An autonomous agent is a system interacting with an environment, sensing it, and acting on it over time following a certain agenda [3]. While there are different classes of agents, in this paper we will be focusing on LLM-based agents which, by combining the capabilities of LLMs and autonomous agents, can achieve a broad range of tasks [1].
@@
In fact, the shift towards more natural conversational interfaces powered by LLMs is transforming the way humans engage with agents, enabling seamless and intuitive interactions. Furthermore, LLM-based agents play an essential role as information fusion intermediaries by intelligently synthesizing data from heterogeneous sources and providing meaningful insights to human users or other AI systems. These LLM-powered agents have now become dominant in the landscape of AI agents, and they are regarded as potential steppingstones towards Artificial General Intelligence (AGI), offering hope for the development of AI agents that can adapt to diverse scenarios [4].
@@
Understanding the potential of these agents and improving them necessitates a deep understanding of their structure. In LLM-based agents, besides the LLM who handles reasoning, there are other components that oversee the execution of tasks, ensure the security of the agent, and handle its memory [4]. But merely pinpointing existing functionalities is insufficient for developers and researchers. To provide a comprehensive overview of these agents, survey [5] proposed a comprehensive framework for their construction, consisting of four main modules which will be discussed thoroughly in the next section. Although it provides valuable insights into the functionality of each module in an agent, it overlooks their delineation from a software architectural perspective. This perspective is crucial for developers to establish a common base architecture to build upon and for researchers to improve. Our analysis of the state-of-the-art LLM-based agents reveals common limitations: the complexity of implementation, resulting in unstructured and ambiguous software architecture; lack of modularity and composability, making components non-reusable by other agent-based solutions; and difficulty in maintainability and introducing improvements to existing agents.
@@
In response to these limitations, we propose the LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF). To the best of our knowledge, our framework is the first to emphasize a clear delineation of each component within an LLM-powered agent and define their interactions within specified boundaries from both architectural and functional perspectives. In addition to traditional components like LLMs and tools, we introduce a new unit within the agent, which we label as the "core-agent". This pioneering entity is further classified into two types – active and passive core-agents – which enhance our understanding of each component’s capabilities and accurately describe the dynamics between modules within the agent. As a result, we alleviate the complexity of the architecture and improve the reusability of the components, promoting a shift from multi-agent systems to multi-core agents.
@@
Throughout our paper, we highlight several advantages offered by the LLM-Agent-UMF, from resolving terminological ambiguities to the introduction of enhancing modules like the security module. To validate the reliability of our framework, we apply it to existing solutions and identify the modules within each agent. This approach enables us to assess the feasibility and requirements for merging one agent with another, ultimately aiming to create an agent with fully enhanced capabilities.
@@
Compared with previous works, the five main contributions of this paper can be summarized as follows:
@@
    We introduce a new terminology, core-agent, as a structural sub-component in LLM-based agents to improve modularity and promote more effective and precise communication among researchers and contributors in the field of agents and LLM technologies.
@@
    We model the internal structure of a core-agent by adapting the framework suggested by [5] that was originally meant to describe the whole agent from an abstract functional perspective.
@@
    We improve our modeling framework by augmenting the core-agent with a security module and introducing new methods within other modules.
@@
    We classify core-agents into active and passive ones, explaining their differences and similarities, and highlighting their unique advantages.
@@
    Finally, we introduce various multi-core agent architectures, emphasizing that the hybrid one-active-many-passive core-agent architecture is the optimal setup for LLM-based agents.
@@ 2 Related Work
We will start off this section by providing a comprehensive overview of the key concepts that serve as the fundamental basis for our work. First, tool-augmented LLMs are a major advancement in Natural Language Processing (NLP) that combine the language understanding and generation capabilities of LLMs with the ability to interface with external tools and Application Programming Interfaces (APIs). For instance, TALM [6] introduces models which can leverage a wide range of functionalities, from information retrieval to task planning and execution.
@@
Figure 1: Framework proposed by survey [5] for LLM-based agents
@@
3 LLM-based Agent Unified Modeling Framework
In order to comprehensively explore LLM-based agents’ capabilities, a rigorous study was conducted across numerous scholarly databases and online digital libraries. These sources include ScienceDirect, Scopus, Springer, Nature, IEEE Xplore, ACM Digital Library, ACL Anthology, NeurIPS proceedings, AAAI library, arXiv, Google Scholar, Semantic Scholar, among others. The analysis was conducted on articles focusing on tool-powered LLMs, LLM-based agents, LLM planning strategies, information fusion, AI ethical concerns, AI privacy considerations, LLM security and guardrails, software modeling best practices, and terminological consistency in the context of AI systems.
In this section, we start by introducing the fundamental component of our framework, the core-agent, distinguishing it from other elements within an LLM-based agent. By examining the existing literature and analyzing their shortcomings, we delineate a unified approach to designing the internal structure of the core-agent. This comprehensive analysis led us to classify core-agents into two categories: active and passive core-agents. Finally, we outline various multi-core agent architectures that highlight the efficiency of LLM-Agent-UMF in modeling multi-core agent systems.
@@ 3.1 Core-Agent: Keystone component of an LLM-based agent
Figure 2: The core-agent as the central component of LLM-based agents
@@
The LLM acts as a cerebral entity covering cognitive tasks such as natural language understanding and comprehensive text generation based on specific context. Indeed, LLMs can communicate with core-agents which in turn can interact with tools or data sources within the system. Considering that an agent can manifest as unimodal or multimodal [18], the latter can be replicated with different task-specific LLMs, as illustrated in Figure 2. This distinction fosters modularity within our LLM-based agent unified modeling framework, making them open for extension and addition of further domain-specific models without having to introduce changes to the overall system architecture, adhering more to the Open-Close principle (OCP) [19, 20].
@@
In LLM-based agents, tools are important assets. They can take various forms such as supplementary systems, software applications or even physical devices [12] that extend and enhance the capabilities of an agent. They span across a spectrum of complexity, ranging from basic API integration to sophisticated auxiliary systems designed for specific tasks. In this context, tools can be considered external if they are independent systems which achieve a complete objective, or internal if they cooperate with the core-agent to achieve tasks in the scope of the agent goal.
@@ 3.2 Modeling the Core-Agent Internal Structure
The human brain exhibits remarkable modularity, with distinct regions and functionalities working in coordination to facilitate cognitive processes, decision-making, and behavior. Inspired by this organized design, our proposed framework aims to emulate the brain’s modular structure thought the incorporation of five internal modules: the planning module, the memory module, the profile module, the action module and the security module (Figure 3). By integrating the latter concepts, the core-agent serves as the central coordinator, controlling the interactions and information flow between different components, as elaborated in the previous section.
Figure 3: Overview of the core-agent internal structure within an LLM-based agent
@@
To provide a robust comparison and underscore the novelty of our framework, it is instructive to examine it alongside existing approaches. A recent survey [5] introduced a framework for LLM-based agents, comprising four key modules: profile, memory, planning, and action modules. For a comprehensive overview of this framework, we direct readers to Section 2. While the survey’s proposed framework treated the LLM-based agent as a whole system without distinguishing a core component from other entities, our adaptation, tailored specifically for the core-agent, necessitates modifying certain definitions and module structures to accommodate the separation between the LLM and the core-agent. Notably, we had to reposition the four key modules to operate under the core-agent within the framework alongside our newly introduced security-oriented module.
@@ 3.2.1 Planning Module
This section will detail the functionalities and characteristics of this module from four critical aspects: process, strategies, techniques, and feedback sources (Figure 4). These aspects derive from the clear separation we have established between the core-agent and the LLM, and between our system and external systems (Figure 3).
Figure 4: Planning module functional perspectives
@@
    This initial phase involves breaking down a complex task into simpler subtasks, thereby establishing a structured hierarchy of intermediate goals. The decomposition of complex tasks can adopt two primary approaches: In the non-iterative decomposition approach, the complex task is broken down into simple subtasks all at once. The planning module defines all subtasks, creating a complete task hierarchy in a single step. This method provides a comprehensive overview of the entire task structure upfront. However, the iterative decomposition approach involves a step-by-step breakdown of the task. In fact, the planning module first defines the initial subtask and goal to reach, establish a proper plan and generates the procedures to undertake. After performing the procedures and completing the plan, the output is considered, and the next subtask is defined. This process is repeated, with each new subtask being planned and executed, contributing to defining the next subtask to achieve. The key advantage of this method is its flexibility and ability to adapt the plan based on the outcomes of each subtask. An exemplary application of the iterative approach is the Decomposition-Alignment-Reasoning Agent (DARA) framework [23]. DARA demonstrates how iterative decomposition can lead to more precise and context-aware planning, particularly for complex tasks with multiple interdependent subtasks.
@@
    This approach involves generating a singular path or sequence of procedures to achieve the goal, adhering to the plan step-by-step without exploring alternatives, thereby providing a straightforward, deterministic approach to planning. Chain of thought (CoT) [24] is an example that outlines such a strategy. Indeed, it uses sequential reasoning as it involves breaking down complex problems into multiple procedures, each built on the previous ones.
@@
    As a matter of fact, the Tree of Thoughts (ToT) and Graph of Thoughts (GoT) [25] are two frameworks that utilize this multi-path approach. They both operate by leveraging an LLM as a thought generator to produce intermediate procedures, that are structured either as a hierarchical tree in case of ToT or as a more complex graph in case of GoT. However, the complexity of managing these structures cannot be solely handled by the LLM. It requires the integration of a specialized software component responsible for orchestrating the process, interacting with the LLM, and organizing the thoughts into the desired structure, whether a tree or a graph. This essential software component is identified as the planning module within the core-agent. The planning module further evaluates different paths within the generated structure and selects an optimal plan [22] based on its assessment.
@@
The planning module follows planning techniques as methodological approaches to form executable plans. These techniques are chosen based on criteria such as the complexity of the task, and the need for contextual comprehension. Our framework presents two primary techniques:
@@
    Within our architectural framework, rule-based methodologies encompass what is commonly referred to in literature as symbolic planners [22]. These techniques proved to be valuable especially in contexts characterized by complex constraints, such as mathematical problem-solving or the generation of plans within highly problematic situations.
@@
Planning without feedback may pose several challenges as feedback plays a crucial role in optimizing the performance of the planning module within the core-agent. For instance, in iterative task decomposition, feedback has an influential impact on the next generated step and enhances the agent’s alignment with the user’s expectations. To effectively address these challenges, the planning module relies on diverse feedback sources. As outlined in Figure 3, the core-agent engages with tools within its system boundaries, as well as entities outside its scope, such as external systems and humans. Consequently, interactions with these components can offer valuable feedback:
@@ 3.2.2 Memory Module
The memory module is responsible for the storage and retrieval of information pertinent to the core-agent’s activities, thereby enhancing the core-agent’s decision-making efficiency and task execution capabilities. In [27] and [5], the memory module in an LLM-based agent was approached from an abstract functional perspective, neglecting its analysis from a software architectural viewpoint. This led to some overlap between the memory module and the other modules defined in the framework suggested by [5]. Consequently, we propose a more comprehensive and well-defined presentation of this module based on three perspectives: Memory Scope, Memory Location, and Memory Format (Figure 5).
Figure 5: Memory module functional perspectives
The framework presented in the survey [5] classified the memory structure into Unified Memory, intended to emulate human short-term memory via in-context learning, and Hybrid Memory, which represents both short-term and long-term memory functionalities. While these names aimed to differentiate types of agent systems, they deviate unnecessarily from the well-established terminology in the field. A more conventional and widely recognized categorization is introduced through the first perspective, Memory Scope, which includes short-term and long-term memory, aligning more with the human memory types as detailed in [28].
As opposite to how it was defined in framework [5], our short-term memory definition diverges to focus more on the impact of the core-agent rather than on the LLM. From a content perspective, various sources of information contribute to short-term memory. As pointed out in paper [27], data can be derived from one trial of a given task or from previous trials of the same task, which is referred to as short-term memory. This data has a narrow scope that focuses on a specific task and primarily communicated to the LLM for the purpose of in-context learning and enhancing the LLM capability with more information related to the task at hand. Conversely, long-term memory refers to the ability to store and recall information over extended periods, beyond the scope of a specific task. This type of memory enables the core-agent to maintain coherence and context over prolonged interactions, learning and adapting from past experiences. Namely, MemoryBank [29] stores all interactions with user in a large symbolic memory and processes experiences into high-level summaries to reflect upon future similar tasks.
@@
Beyond focusing on the scope of the data, inside trial and across trials, as stated in [27], we emphasize the location of the memory because it is more relevant from a software architecture point of view. Hence, we introduce the second perspective, Memory Location, which encompass two categories: Embedded Memory, internal to the core-agent boundaries, and Memory Extension, outside the core-agent boundaries, yet still within the boundaries of the agent system as depicted in Figure 3.